Consider a particle of mass m on x-y plane.The particle is moved by a force F~ =xy ˆi + y jˆ fromthe origin (0, 0) to a point (1, 1) along a straight

Line joining the two pints. Find the workdone

By the force.



1
Expert's answer
2021-02-08T18:37:46-0500

"F=xy\hat{i}+y\hat{j}"

"W=\int_0^1dw=\int_0^1Fds"

"W=\int_0^1dw=\int_0^1Fdx\hat{i}+\int_0^1Fdy\hat{j}"

"\text{moving from (0, 0) to point (1, 1) in a straight line}"

"\text{is described by the equation } y=x"

"dy=dx"

"W=\int_0^1xydx\hat{i}+\int_0^1ydy\hat{j}="

"\int_0^1x^2dx\hat{i}+\int_0^1xdx\hat{j}=(\frac{x^3}{3}+\frac{x^2}{2})|_0^1=\frac{5}{6}"


Answer:"\frac{5}{6}N"